OpenStudy (anonymous):

Solve the DE using method of undetermined coefficients y"'-6y"+11y'-6y=2xe^-x

OpenStudy (anonymous):

hi you can start with finding you Yc y"'-6y"+11y'-6y=2xe^-x y"'-6y"+11y'-6y=0 m^3 -6m^2 +11m -6=0 m=1,2,3 therefore Yc=C1e^x +C2 e^2x +C3 e^3x now find Yp by using Yp=(Ax+B)e^-x Yp'=___? Yp''=____? Yp'''=_____? then sub them to the original equation

OpenStudy (anonymous):

@ briangtt